OpenStudy (anonymous):

PLEASE HELP? Given: The corrdinates of triangle PQR are P(0,0) , Q(2a,0) , and R(2b,2c). Prove: The line containing the midpoints of two sides of a triangle is parrallel to the third side. As part of the proof, find the midpoint of PR?

OpenStudy (anonymous):



OpenStudy (anonymous):

the midpoint between (x1,y1) and (x2,y2) is\[\left( \frac{x_{1}+x_{2}}{2},\frac{y_{1}+y_{2}}{2} \right)\]

OpenStudy (anonymous):

Still kind of lost? /:

OpenStudy (anonymous):

find the midpoint of PR first